### v1

# Scheduled Curiosity-Deep Dyna-Q: Efficient Exploration for Dialog Policy Learning

Abstract
Training task-oriented dialog agents based on reinforcement learning is time-consuming and requires a large number of interactions with real users. How to grasp dialog policy within limited dialog experiences remains an obstacle that makes the agent training process less efficient. In addition, most previous frameworks start training by randomly choosing training samples, which differs from the human learning method and hurts the efficiency and stability of training. Therefore, we propose Scheduled Curiosity-Deep Dyna-Q (SC-DDQ), a curiosity-driven curriculum learning framework based on a state-of-the-art model-based reinforcement learning dialog model, Deep Dyna-Q (DDQ). Furthermore, we designed learning schedules for SC-DDQ and DDQ, respectively, following two opposite training strategies: classic curriculum learning and its reverse version. Our results show that by introducing scheduled learning and curiosity, the new framework leads to a significant improvement over the DDQ and Deep Q-learning(DQN). Surprisingly, we found that traditional curriculum learning was not always effective. Specifically, according to the experimental results, the easy-first and difficult-first strategies are more suitable for SC-DDQ and DDQ. To analyze our results, we adopted the entropy of sampled actions to depict action exploration and found that training strategies with high entropy in the first stage and low entropy in the last stage lead to better performance.
Keywords Dialog management  $\cdot$ reinforcement learning  $\cdot$ Deep Dyna-Q  $\cdot$ curiosity  $\cdot$ curriculum learning
1 Introduction
Since human-computer interaction and natural language processing are in high demand in industry and daily life, and the task-oriented dialog system has become a hot topic and deserves further study and research. Dialog policy model is used to select the best action at each step of a dialog. In a task-oriented dialog system, the goal of the dialog is to complete a specific task, such as booking a movie. The dialog policy should select actions to efficiently achieve the goal. Dialog policy learning is often formulated as a reinforcement learning (RL) problem [1, 2], in which a dialog agent executes an action based on the observed state and receives a reward from the environment acted by real users. However, optimizing RL agents requires a large number of interactions with the environment, which is an expensive and time-consuming process. Therefore, how to grasp dialog policy efficiently within limited interactions remains a fundamental question for RL [3, 4, 5, 6, 7]. A common approach is to encourage the agent to explore the environment as sufficiently as possible through limited interactions. In contrast to supervised and unsupervised learning, reinforcement learning does not contain a large amount of training data with or without labels prepared in advance, but rather generates experience through the agent’s interactions with the environment and continually optimizes the agent based on feedback from the environment. Therefore, exploration of the environment is crucial for RL, and extensive research has been conducted on promoting agents to explore the environment. Count-based exploration is an intuitive approach in which each state observed by the agent is recorded and the agent is encouraged to learn states with fewer occurrences. The most evident drawback of this approach is its inability to handle complex and high-dimensional data. To address this, Bellemare et al. engineered a model that acquires the distribution of visited environment states to measure the novelty of states and encourage agents to observe unfamiliar states [8]. In addition, to deal with high-dimensional data, Tang et al. suggested mapping high-dimensional states to hash codes [9], and Liu et al. introduced embedding networks to encode the state space [10]. However, a common flaw in such approaches is that they do not lead agents to explore states that have never existed before. Another method is to model the intrinsic motivation of the agents. One ingenious strategy is to utilize curiosity to facilitate state exploration. Pathak et al. [11] proposed an Intrinsic Curiosity Module (ICM) to model the state prediction error, which depicts the uncertainty and improvement in RL as a curiosity reward to encourage unfamiliar dialog states. This research makes some simplifications and adjustments to ICM to make it more suitable for task-oriented dialog systems to guide unfamiliar dialog state exploration. This model leveraged the predictability of a state to depict its novelty. If the next state can be accurately predicted based on the current state and the upcoming action to be performed, it means that the action cannot lead the agent to see unfamiliar states, and it corresponds to a less curiosity reward. By contrast, if the next state is difficult to predict, the corresponding action receives a larger curiosity reward. The agent executes the action with the largest sum of curiosity rewards and rewards from the environment. Furthermore, the RL-based dialog agent faces another challenge. Generally, an episode of dialog based on RL is opened up by randomly selecting a user goal from the entire training dataset. Because the subsequent conversation revolves around the drawn goal until it is achieved successfully, it has an important influence on the subsequent training process. Random sampling neglects the way humans acquire knowledge and skills, where they focus on relatively easy materials before harder ones, and hurts the efficiency and stability of the training process. To overcome this problem, Bengio et al. [12] proposed curriculum learning (CL), which presents relatively easy or simple examples at an early stage, inspired by human learning habits. Many NLP tasks have been improved by using a typical CL [13, 14, 15, 16, 17, 18, 19, 20]. These learning methods employed different evaluations of training examples and approaches to adjust the training steps but shared the easy-first strategy. However, in some studies, the reverse version of CL was tested and achieved the best performance among various training scheduler designs [18, 21, 22]. The effectiveness and application of easy-to-hard and hard-to-easy strategies are still worth exploring [23]. Chang et al. [21] stated that the difficult-first strategy is more suitable for cleaner datasets, whereas the classic CL is beneficial for acquiring policies through noisy scenarios and leads to faster convergence. Furthermore, when tasks are difficult for an agent to complete, earlier presentation of easier samples is preferable for an effective training process. However, the exact impact of scheduled training on agent behavior and policy optimization remains unclear. This study aims to improve the performance of a task-oriented dialog system by providing sufficient environment exploration and a training strategy that matches human learning habits. Therefore, we propose Scheduled Curiosity-Deep Dyna-Q (SC-DDQ) which combines CL and a curiosity reward with dialog policy optimization based on Deep Dyna-Q (DDQ), where DDQ is a state-of-the-art model-based dialog system [7], and design a curriculum based on two opposite learning strategies, which are found to be optimal in different scenarios. Compared to count-based exploration approaches [8, 9, 10], and variants of DDQ, SC-DDQ for the first time adapts the curiosity model which models the agent’s intrinsic motivation, and introduces it into DDQ to motivate the agent to explore the environment based on its intrinsic motivation. In addition, SC-DDQ develops criteria for classifying the difficulty of training samples based on the capabilities of the dialog agent and introduces CL into the training of the dialog agent. Because the specific application scenarios of easy-to-difficult and difficult-to-easy learning strategies are not yet conclusive, in the experimental section, we carried out four combinations of experiments based on the presence or absence of the curiosity model and the trend of the difficulty of the task, aiming to explore the effects of agent intrinsic motivation and learning strategies on the completion of the dialog tasks. Preferring more difficult user goals is effective when experiments are conducted without curiosity reward while focusing on easier tasks is found to be an optimal training strategy if the dialog agent selects its action according to both external and internal rewards. To determine and analyze the influences of scheduled training on agent behavior, the entropy of agent action sampling in each training stage, measuring the agent action exploration, was employed to characterize the policy strategy. It is worth noting that entropy is designed to analyze the experimental data, not part of the algorithm itself. We point out that a higher entropy at an earlier stage yields better performance. In other words, at the beginning of learning, relatively uniform action sampling, namely more sufficient action space exploration, leads to a higher task success rate. According to the experimental results, a key factor in facilitating dialog task completion was encouraging action exploration during the early training stage. Compared with [3, 4, 5, 8, 9, 10], SC-DDQ is more scalable. First, this study makes some simplifications to ICM which is the original curiosity model. It performs feature extraction on the state to filter out parts of the state that are unaffected by action execution. Because almost all elements contained in the representation of the dialog state in the context of movie ticket booking scenarios are manipulated by action execution, feature extraction is discarded in this research, which should be borrowed for other scenarios whose states do not contain too many interfering parts. Second, the introduction of CL is mainly an adaptation of the training sequence; therefore, it can be easily extended to algorithms other than dialog systems. Furthermore, our experimental results provide evidence that different training scenarios can potentially benefit from both classical CL and its inverse version. This result presents new ideas for subsequent research on CL. Therefore, the SC-DDQ framework is scalable to real-world applications such as cyber defense. Moradi et al. leveraged reinforcement learning to acquire an attack strategy and allocate defense resources to protect smart electrical power grids[24, 25]. The main idea of SC-DDQ can be adapted to these two cyber defense frameworks. For example, the curiosity reward can be modeled as the prediction error of the defense resource allocation status, and the order of the training tasks presented to the agent can be adjusted according to their defense difficulties. In summary, our main contributions in this paper are three-fold:
  1. (1)
    We propose a curiosity-driven scheduled framework extended on the Deep Dyna-Q (DDQ) model to improve the performance and learning efficiency of task-oriented dialog systems. To the best of our knowledge, this is the first adaptation of the curiosity model, integrating it into a task-oriented dialog system. This result is presented in Section 5.1.
  2. (2)
    We designed learning curricula based on opposite training strategies and presented both benefits to policy grasping under different reward settings. Compared with common applications of CL, this research is not limited to an intuitive easy-to-difficult training strategy, and for the first time, opposite training strategies are implemented in different settings of the same framework. This contribution is presented in Section 5.2.
  3. (3)
    We adopt the entropy of agent action sampling to depict the behavioral characteristics of the agent, and point out that guiding the agent to attempt various actions in the early phase of training facilitates policy optimization, which is explained in detail in Section 5.3.
2 Related work
2.1 Deep Dyna-Q
Our research is based on the Deep Dyna-Q (DDQ) model, a classic model-based RL dialog system that integrates planning to improve the task completion rate within limited interactions [7]. The framework is illustrated in Fig. 2. It consists of three processes: (1) direct reinforcement learning, in which the policy model learns from real dialog experience; (2) planning, in which the world model is applied to generate a simulated experience to improve the dialog policy model; and (3) world model learning, in which the world model is improved through real experience.
2.1.1 Direct Reinforcement Learning
Dialog policy learning can typically be formulated as a Markov Decision Process (MDP). An episode of task-oriented dialog can be regarded as a set of tuples. In each dialog turn, the policy agent observes the dialog state $s$ and samples action $a$. The selection of action $a$ is based on $\varepsilon$-greed policy, where the action is chosen randomly with probability $\varepsilon$ or to maximize the action-value function $Q(s,a;\theta)$. The action-value function is accomplished by a Multi-Layer Perceptron (MLP) with parameters $\theta_{Q}$. After executing this action, the agent receives a reward $r$ from the environment and observes the user response $a^{u}$. The dialog state is then updated to the next state $s^{\prime}$. The tuple $(s,\ a,\ r,\ a^{u},\ s^{\prime})$ can be viewed as a piece of experience and stored in either a real or a simulated experience replay buffer. The action-value function $Q(\cdot)$ is improved via Deep Q-network (DQN) [6]. An episode of dialog is launched by sampling a user goal from a goal set by the user simulator. Each user goal is defined as G = (inform_slots, request_slots), where inform_slots is a set of constraints and request_slots is a set of requests. For the movie-ticket booking task, typical information slots involve items such as movie names or number of people. Requests can be theater or start time.
Figure 1: The framework of DDQ
Figure 1: The framework of DDQ
Figure 2: The structure of world model
2.1.2 Planning and World Model Learning
In the planning process, the world model, which simulates the environment, interacts with the policy agent and generates simulated experiences stored in a simulated replay buffer. In each planning turn, state $s$ and agent action $a$ are viewed as inputs to the world model. The world model then generates its response $a^{u}$, the corresponding reward $r$, and a binary variable $t$ indicating whether this episode is over, as shown in Fig. 2. Both direct reinforcement learning and planning were accomplished using the same DQN algorithm, training on real and simulated experience. The world model was implemented as an MLP and tuned based on real experiences during the world model learning process.
2.2 Curriculum Learning
Curriculum learning (CL) is a training strategy in which the agent initially concentrates on relatively easy data and progressively moves to harder training samples. This strategy is inspired by human study characteristics and is beneficial for accelerating the training convergence speed and enhancing the agent performance. CL is more implementable and compatible than multitudinous and complicated frameworks and models, thereby making it more universally applicable. In addition, CL makes the best use of the existing data, which saves resources and time for RL-based tasks. Typical CL have been used extensively in natural language processing. Liu et al. [17] employed the mastering level of current tasks as the difficulty criterion and guided the agent to grasp easier tasks first. Zhao et al. [19] utilized the total number of inform and request slots for tasks to identify tasks, and the task completion rate to adjust the difficulty level of the next task. However, the easy-to-hard learning strategy is not always helpful. In some studies, the opposite version of CL, focusing on harder examples, brings more significant improvements to the system. Zhao et al. [18] proposed an opposite CL strategy, in which the agent was trained on a general training set and gradually moved to subsets. Moreover, Hacohen et al. [22] demonstrated that self-paced learning, a well-known variant of CL, hurts the agent’s performance. Chang et al. [21] stated that CL is more suitable for training scenarios with significant noise. By contrast, the training process of cleaner scenarios would be more efficient using the opposite CL strategy. As the effectiveness and applicability of difficult-first and easy-first strategies are still open to debate, we utilized the difficulty of user goals as a measurement of dialog complexity and proposed opposing learning schedules based on CL and its reverse version. Applying different reward functions yields a performance boost for the dialog agents based on the two learning schedules.
2.3 Curiosity Reward
Curiosity is modeled by state error prediction to promote the agent to attempt unconversant states [26, 11]. In RL applications [27, 28, 29, 30, 31], an agent samples an action and receives a reward after observing the current state. Subsequently, the entire dialog environment updates to the next state. If a new state can be predicted accurately before selecting the following action, it is viewed as a familiar state for the agent. However, a state that is difficult to forecast is more informative for the agent. In this study, we integrate the curiosity model with dialog policy optimization to guide agents to explore unfamiliar environment states. In this research, we simplified and adjusted the Intrinsic Curiosity Module (ICM) [11]. ICM is engineered to play video games, such as VizDoom and Super Mario Bros, and it observes screenshots of the game screen as inputs. Because the game screen often contains many disturbances that are not controlled by the action, such as the change of scenery in the background, ICM designs an inverse model to filter out these disturbing factors and thus performs feature extraction on the states. Unlike the application scenarios of ICM, the states in this research are affected by the execution of actions and contain almost no interference factors. Therefore, the curiosity model used in this research removed the inverse model. In addition, ICM calculates the prediction error as the curiosity reward between the next state and the predicted one after the agent executes the action and observes the true next state. In this research, the agent does not execute the feasible actions individually to observe the corresponding next state and calculate the curiosity reward; instead, the agent takes both the curiosity reward and the predicted next state as outputs for training. This eliminates the need for the agent to repeatedly step back, but the disadvantage is that at the early stage of training, the curiosity reward output from our curiosity model may not be accurate enough. Previous studies introduced curiosity-based exploration from a video game scenario into an RL-based dialog system and achieve performance breakthroughs. Wang et al. [32] investigated different exploration strategies, including ICM, in task-completion dialog policy learning and showed improvements. Doering et al. [33] developed a shopkeeper robot whose verbal interactions with customers are guided by curiosity, and it is significantly human-like, compared to non-curious robots. These studies provide evidence for the potential effectiveness of the curiosity model in dialog systems. Therefore, we made some adjustments to the ICM, integrated it into the dialog system, and believe in its effectiveness.
3 Proposed methods
3.1 Overview
The common training strategy for dialog agents is to randomly expose the agent to tasks with different difficulties, where the training efficiency is low and the agent’s performance can be hurt. In addition, the dialog environment acted by real users is too complicated to be fully explored. Therefore, we propose Scheduled Curiosity-Deep Dyna-Q (SC-DDQ), a novel and practical framework applying a curiosity strategy to joint curriculum learning and RL-based policy learning for task-oriented dialog systems, implemented based on Deep Dyna-Q to improve the dialog agent performance.
Figure 3: The framework of Scheduled Curiosity-Deep Dyna-Q
The proposed framework is illustrated in Fig. 3. The SC-DDQ framework consists of three modules: (1) an off-line task classifier for dividing user goals into three complexity levels; (2) a curiosity policy agent for selecting the next action beneficial for completing the task and exploring state by using the current dialog state; and (3) a world model for simulating a user to generate actions and rewards based on real user behaviors. The optimization of SC-DDQ comprises five steps: (1) warm starting: a hand-crafted dialog policy is employed to generate experiences; (2) direct reinforcement learning guided by curiosity: the agent guided by the curiosity model interacts with real users and generates real experiences. They are stored in the real experience replay buffer and used to improve dialog policy; (3) world model learning: the world model is refined on real experiences; (4) planning guided by curiosity: the agent guided by the curiosity model conducts interactions with the world model, where the simulated experiences are collected and employed to optimize dialog policy; and (5) curiosity model training: the curiosity model is optimized using both real and estimated experiences. In addition, the proposed framework can be equipped with different policy models and curriculum schedules. Four combinations are implemented in this research, as shown in Table 1: (1) policy agent without curiosity equipped with the easy-first strategy (EFS) curriculum; (2) policy agent without curiosity equipped with the difficult-first strategy (DFS) curriculum; (3) curiosity-combined policy agent equipped with the EFS curriculum; and (4) curiosity-combined policy agent equipped with the DFS curriculum.
Table 1: Four combinations
Method Curiosity Schedule
DDQ No Random
C-DDQ Yes Random
S-DDQ_EFS No EFS
S-DDQ_DFS No DFS
SC-DDQ_EFS Yes EFS
SC-DDQ_DFS Yes DFS
In this research, two opposite types of curriculum training schedules are proposed, both of which enable the RL dialog agent to achieve better performance under certain conditions. To reveal the influence of the training curriculum on the performance of the agent, the response of the agent in each turn of the dialog was recorded. These results are presented in Section 5.2. The structure of the main parts of SC-DDQ is described in the subsequent subsection. The pseudo-code for the iterative SC-DDQ algorithm is shown in Algorithm 1. Line 1 describes the first step in the optimization, where a rule-based dialog is conducted. Lines 2-12 show that the off-line task classifier divides user goals into three levels. Lines 14-17 illustrate that before launching an episode of dialog a user goal that fits the curriculum is sampled. The second step of optimization is described in lines 18-24, where the curiosity policy model outputs Q-values and curiosity rewards and selects the best feasible action based on their sum. Line 24 describes the update of the DQN based on real experiences. The third step of optimization corresponds to line 25 where the world model is updated using real experiences. Line 26 illustrates the fourth step, where the world model, replacing the user simulator, interacts with the agent and generates simulated experiences. Line 27 shows that the DQN is optimized by simulated experiences. The last step of the optimization is shown in line 28, where the curiosity model is updated using both real and simulated experiences.
3.2 Task Classifier
User goal:
"request_slots": {
"ticket": "UNK"
"diaact": "request",
"inform slots": {
"city": "Seattle",
"number of people": "2",
"theater": "amc pacific place 11 theater",
"start time": "10:00 pm",
"date": "tomorrow",
"movie name": "race"
Dialog conducted by rule-based agent:
Turn 0 usr:
I want tickets for race at 10:00 pm.
Turn 1 sys:
What movie are you interested in?
Turn 2 usr:
I want to watch race.
Turn 3 sys:
What time would you like to see it?
Turn 4 usr:
I want to watch at 10:00 pm.
Turn 5 sys:
Which city would you like?
Turn 6 usr:
I want to watch at Seattle.
Turn 7 sys:
What date would you like to watch it?
Turn 8 usr:
I want to set it up tomorrow
Turn 9 sys:
Which theater would you like?
Turn 10 usr:
I want to watch at amc pacific place 11 theater.
Turn 11 sys:
How many tickets do you need?
Turn 12 usr:
I want 2 tickets please!
Turn 13 sys:
Great - I was able to purchase 2 tickets for you to see race tomorrow at amc pacific place 11 theater in Seattle at 10:00 pm.
Turn 14 usr:
Thank you.
Turn 15 sys:
Thank you.
Figure 4: A success task completed by rule-based agent
Table 2: The data annotation schema
Intent
request, inform, deny, confirm_question,
confirm_answer, greeting, closing, not_sure,
multiple_choice, thanks, welcome
Slot
city, closing, date, distanceconstraints,
greeting, moviename, numberofpeople, price,
starttime, state, taskcomplete, theater,
theater_chain, ticket, video_format, zip
[…]
Easy goal
Request slots Inform slots
1 Ticket UNK 1 City Seattle
2 Number of people 2
3 Theater Royal theater
4 Start time 10:00 pm
5 Date Tomorrow
6 Movie name Race
(a) An easy goal
Middle goal
Request slots Inform slots
1 Ticket UNK 1 Number of people 3
2 Theater UNK 2 Date Tonight
3 Start time UNK 3 Movie name Zootopia
(b) A middle goal
Difficult goal
Request slots Inform slots
1 Ticket UNK 1 City Los Angeles
2 Date UNK 2 Number of people 1
3 Theater UNK 3 Movie name Deadpool
4 Start time UNK
(c) A difficult goal
Figure 5: User goals in different difficulty level. UNK means that the corresponding slot is unknown.
3.3 Curiosity Policy Model
The proposed curiosity policy model is a combination of the DQN-based agent [6] and the curiosity model, as shown in Fig. 6. The DQN-based agent $Q(s_{t},a_{t};\theta_{Q})$ is used to interact with users, and the curiosity model $C(s_{t},a_{t};\theta_{C})$ is used to generate a curiosity value to assist the agent in selecting an action that leads to an unfamiliar state.
Figure 6: The structure of the curiosity policy model
During direct reinforcement learning, in step $t$, the agent observes state $s_{t}$ and chooses an action $a_{t}$ to carry into the next dialog turn. Traditionally, the action $a_{t}$ is chosen according to $\epsilon$greedy, where we choose a random action with probability $\epsilon$, otherwise the action is chosen following the greedy policy $a_{t}\ =\text{argmax}_{a^{\prime}}Q\left(s_{t},\ a^{\prime};\theta_{Q}\right)$. The $a^{\prime}$ represents feasible agent actions, and $Q\left(s,\ a;\theta_{Q}\right)$ is the approximated value function, implemented as a Multi-Layer Perceptron (MLP) parameterized by $\theta_{Q}$. To encourage the exploration of unfamiliar states, we introduce the curiosity value generated by the curiosity model in this step. The curiosity model was implemented using MLP $C\left(s,\ a;\theta_{C}\right)$. As shown in Fig. 6, this model takes the current state $s_{t}$ and candidate actions $a^{\prime}$ as inputs, and the predicted next state $s_{t+1}^{\prime}$ and the curiosity value $c_{a^{\prime}}$ as outputs. A larger curiosity value indicates a larger difference between the real and predicted next state. For an agent, a state that cannot be predicted accurately is unfamiliar and informative. Therefore, the action maximizing the addition of the Q value and curiosity value is sampled, which can be formulated as Equation 1:
$a_{t}=argmax_{a^{\prime}}\left[Q\left(s_{t},\ a^{\prime};\theta_{Q}\right)+c_{a^{\prime}}\right]$ (1)
Subsequently, the agent receives a reward $r_{t}$ from the environment, observes the next user action $a^{u}$, and updates to the next state $s_{t+1}$ until it reaches the end of the dialog. Experience $(s_{t},\ a_{t},\ r_{t},\ a^{u},\ s_{t+1})$ is stored in the real replay buffer $D^{u}$. We improve the Q value function $Q\left(s,\ a;\theta_{Q}\right)$ using real experiences from $D^{u}$ via minibatch SGD. It should be noted that the weight of the curiosity value cannot be adjusted for the time being. In other words, curiosity-based exploration cannot be adjusted adaptively in this study. In the planning and world model learning process, as described in the related work section, the world model is applied to interact with the agent and generate simulated experiences, which can also be used to improve the dialog policy. The planning process was similar to that of direct reinforcement learning. In this process, the environment is acted by the world model accomplished by an MLP $M\left(s,\ a;\theta_{M}\right)$ as shown in Fig. 2. In each planning turn, the world model takes the current state $s_{t}$ and the agent action $a_{t}$ as inputs. The outputs are user action $a^{u}$, reward $r_{t}$, and a binary variable $t$ indicating whether this episode is over or not. Dialog experiences generated during planning are stored in the simulated replay buffer $D^{u}$. The proposed framework follows the DDQ structure [7], which has two experience replay buffers $D^{u}$ and $D^{s}$ for storing real and simulated experiences, respectively. Because the role of the world model is to imitate real users, real experiences from $D^{u}$ were employed for its optimization via minibatch SGD. In the curiosity model training process, the curiosity model accomplished by MLP $C(s,\ a,\ \theta_{C})$ is refined via minibatch SGD using experiences from both real and simulated replay buffers, namely $D^{u}$ and $D^{s}$. Its inputs are the current state $s_{t}$ and candidate actions $a^{\prime}$, whereas its outputs are the predicted next state $s_{t+1}^{\prime}$ and curiosity value $c_{a^{\prime}}$. The labels of the predicted next state $s_{t+1}^{\prime}$ and the curiosity value $c_{a^{\prime}}$ are the real next state $s_{t+1}$ and state prediction error, respectively. To define the error of state prediction, we encode real and predicted next states $s_{t+1}$ and $s_{t+1}^{\prime}$ into vectors $\phi(s_{t+1})$ and $\phi(s_{t+1}^{\prime})$ accomplished by one-hot encoding. The state prediction error is expressed as $\ \left\|\phi\left(s_{t+1}\right)-\ \phi\left(s_{t+1}^{\prime}\right)\right\|^{2}$.
Input: Dialog user goal set $G_{\text{total}}=\{g_{1},\ldots,g_{n_{\text{total\_user\_goals}}}\}$, $N$
Result: $Q(s,a;\theta_{Q}),M(s,a;\theta_{M}),C(s,a;\theta_{C})$
1 The rule-based agent generates real experiences and stores them into $D^{u}$;
[…]
3       $G_{level}\leftarrow\emptyset$;
4
5 end forall
6for $i\leftarrow 1$ to $n_{\textrm{total\_user\_goals}}$ do
7       switch $N_{\text{request\_slot}}(g_{i})$ do
8             case 1 do $level\leftarrow$ easy;
9              case 2,3 do $level\leftarrow$ middle;
10              otherwise do $level\leftarrow$ difficult;
11
12       end switch
13      Add $g_{i}$ into $G_{level}$;
14
15 end for
16for $n\leftarrow 1$ to $N$ do
17       Determine $level$;
18       Sample a goal $g$ from $G_{level}$;
19       Set the initial state $s_{0}$ and the user’s first action $a_{0}$;
20       $t\leftarrow 1$;
21       while dialog not terminated do
22             DQN $Q(s,a;\theta_{Q})$ generates Q-value;
23             Curiosity model $C(s,a;\theta_{C})$ generates curiosity value $c_{a^{\prime}}$ and the predicted next state $s_{t+1}^{\prime}$;
24             Agent executes action $a=\arg\max_{a^{\prime}}[Q(s_{t},a^{\prime};\theta_{Q})+c_{a^{\prime}}]$, observes the next user action $a^{u}$, receives reward $r_{t}$ and updates to the next dialog state $s_{t+1}$ ;
25             Stores the real experience $(s_{t},a_{t},r_{t},a^{u},s_{t+1})$ to $D^{u}$;
26
27       end while
28      Update $\theta_{Q}$ using real experiences from $D^{u}$ via minibatch SGD;
29       Update $\theta_{M}$ via minibatch SGD on real experiences from $D^{u}$;
30       World model interacts with the curiosity policy agent and generates simulated experiences stored to $D^{s}$;
31       Update $\theta_{Q}$ using simulated experiences from $D^{s}$ via minibatch SGD;
32       Update $\theta_{C}$ using real and simulated experiences via minibatch SGD;
33
34 end for
Algorithm 1 Scheduled Curiosity-Deep Dyna-Q policy learning
In addition, in each training stage, the actions selected by the agent are counted, and sampled action distributions are generated for each agent. The entropy of the sampled agent actions was calculated to reveal the impact of the difficulty level of tasks on the behavior of the agent. It is worth emphasizing that entropy is employed only for the analysis of experimental results and is not part of the algorithm itself.
4 Experimental Setup
4.1 Datasets
The raw conversational data in movie-tickets booking dataset [37]  [^1] https://github.com/xiul-msr/e2e_dialog_challenge is collected via Amazon Mechanical Turk. The dataset was manually labeled based on a schema established by domain experts. The annotation schema has 11 intents and 16 slots, as shown in Table 2. The original dataset contained up to four request slots for user goals. To thoroughly explore the effects of the curriculum on dialog policy, we supplemented nine user goals containing five request slots. Thus, in total, 137 user goals were pre-generated for the movie-ticket booking scenario. As shown in Table 3, there were 61 goals with one request slot, 16 goals with two request slots, 17 goals with three request slots, 34 goals with four request slots, and nine goals with five request slots. Goals with only one request slot are easy. Then, goals with two or three slots were classified as middle. Finally, goals with four or five slots were classified as difficult. A total of 991 movies were available. The information for each movie contains the movie name, date, start time, theater, etc.. There are 29 feasible actions for the agent and 35 feasible actions for the real user and the world model.
Table 3: User goals with different number of request_slot. $N_{\textrm{request\_slot}}$ is the number of request slots, $N_{\textrm{goal}}$ is the number of corresponding goals, and $N_{\textrm{gset}}$ is the total number of goals in the corresponding set.
Goals set $N_{\textrm{request\_slot}}$ $N_{\textrm{goal}}$ $N_{\textrm{gset}}$
Easy 1 61 61
Middle 2 16 33
3 17
Difficult 4 34 43
5 9
4.2 User Simulator
In this research, an accessible user simulator [38] was applied to evaluate the SC-DDQ model to ensure that the training process was affordable and practical. During training, the user simulator offers rewards and simulated user responses according to hand-crafted rules to the dialog agent. An episode of dialog is judged to be successful when the agent provides a suitable movie ticket that satisfies all constraints from the user. The reward rules adopted by the user simulator are as follows: (1) in each turn, a reward of $-1$ is fed back to encourage a shorter dialog; (2) at the end of the dialog, a reward of $2L$ is provided for success or a reward of $-L$ is provided for failure, where $L$ is the maximum number of turns in each episode and is set to 40 in this research.
4.3 Training Schedules
Typical CL suggests presenting relatively easier samples before harder ones, namely the easy-first strategy (EFS). This study employs both the classic CL and its reverse version (difficult-first strategy, DFS) and designs several schedules: EMD, EDD, EED, DME, DEE, and DDM. Following the original DDQ setting, the entire training process contains 300 epochs [7], that were divided into four stages as evenly as possible. The first to third stages, each consisting of 70 epochs, use the corresponding goals of the schedule condition. For example, in the EMD schedule, easy goals are used from 0 to 69 epochs, middle goals from 70 to 139 and difficult goals from 140 to 209 epochs. In the last stage, 210-299 epochs, the goals were uniformly sampled from all goals.
4.4 Examined Methods
We analyzed the impacts of the blue easy-first strategy (EFS) and difficult-first strategy (DFS) by comparing several task-oriented dialog agents that employed variations of Algorithm 1.
DQN
: A task-completion dialog agent learned by standard DQN, implemented by direct reinforcement learning without a curiosity model [6].
DDQ
: A state-of-the-art task-oriented dialog agent as described in the related work section [7].
S-DDQ
: A DDQ agent trained following specific schedules.
C-DDQ
: A DDQ agent equipped with a curiosity model but without curriculum learning.
SC-DDQ
: A DDQ agent equipped with a curiosity model and trained following specific schedules.
4.5 Implementation Details
The DQN $Q(s,a;\theta_{Q})$, curiosity model $C(s,a;\theta_{C})$, and world model $M(s,a;\theta_{M})$ were all accomplished using MLP with tanh activations. Following the original DDQ settings [7], the DQN contains one hidden layer with 80 hidden nodes. The world model had two shared hidden layers and three task-specific hidden layers, each with 80 nodes. The curiosity model contained two hidden layers and two task-specific hidden layers, each with 80 nodes. $\varepsilon$-greedy algorithm was adopted for exploration. Each network was optimized using RMSProp. The batch size was 16, and the discount factor was 0.9. The sizes of the real and simulated replay buffers, $D^{u}$ and $D^{s}$, were set to 5000. The target network is updated at the end of each training epoch. The maximum length of the simulated dialog was 40. When a dialog exceeded the turn limit, it was judged as a failure. In addition, to increase training efficiency, we utilized Reply Buffer Spiking (RBS) [3] and pre-filled the real experience replay buffer $D^{u}$ in the initial training stage with a set of real dialog experiences generated by a rule-based agent [7].
5 Evaluation
The task success rates and the average number of turns are presented in Tables 4 and 5, respectively. Each run was tested on 50 episodes of dialogs. The user goals for testing were sampled from the corresponding goals of the schedule conditions. Three main results are presented in the following subsections.
Table 4: Success rate results. The stage 1 (S1), stage 2 (S2), stage 3 (S3), and stage 4 (S4) shows the results at epoch 70, 140, 210, and 300.
Strategy Method Schedule S1 S2 S3 S4
Random DQN - $0.38$ $0.58$ $0.66$ $0.74$
Random DDQ - $0.64$ $0.60$ $0.88$ $0.86$
Random C-DDQ - $0.62$ $0.88$ $0.90$ $0.90$
EFS S-DDQ EMD $1.00$ $0.58$ $0.00$ $0.74$
EFS S-DDQ EDD $1.00$ $0.00$ $0.56$ $0.88$
EFS S-DDQ EED $1.00$ $1.00$ $0.00$ $0.62$
EFS SC-DDQ EMD $0.98$ $0.58$ $0.64$ 0.92
EFS SC-DDQ EDD $0.98$ $0.54$ $0.66$ $0.88$
EFS SC-DDQ EED $0.98$ $1.00$ $0.66$ $0.88$
DFS S-DDQ DME $0.56$ $0.74$ $1.00$ $0.88$
DFS S-DDQ DEE $0.56$ $1.00$ $1.00$ 0.94
DFS S-DDQ DDM $0.56$ $0.70$ $0.82$ 0.94
DFS SC-DDQ DME $0.00$ $0.10$ $1.00$ $0.56$
DFS SC-DDQ DEE $0.00$ $1.00$ $1.00$ $0.60$
DFS SC-DDQ DDM $0.00$ $0.00$ $0.40$ $0.78$
Table 5: Average turn results. The stage 1 (S1), stage 2 (S2), stage 3 (S3), and stage 4 (S4) shows the results at epoch 70, 140, 210, and 300.
Strategy Method Schedule S1 S2 S3 S4
Random DQN - $31.28$ $23.36$ $25.48$ $23.32$
Random DDQ - $23.12$ $22.36$ $15.44$ $18.20$
Random C-DDQ - $24.57$ $22.73$ $25.32$ $16.68$
EFS S-DDQ EMD $24.48$ $28.00$ $39.09$ $24.88$
EFS S-DDQ EDD $24.48$ $37.40$ $29.12$ $19.56$
EFS S-DDQ EED $24.48$ $17.53$ $39.72$ $24.56$
EFS SC-DDQ EMD $17.52$ $23.72$ $18.12$ 13.36
EFS SC-DDQ EDD $17.52$ $31.72$ $21.44$ $15.64$
EFS SC-DDQ EED $17.52$ $16.73$ $22.88$ $22.20$
DFS S-DDQ DME $23.04$ $17.52$ $16.81$ $18.88$
DFS S-DDQ DEE $23.04$ $19.08$ $17.48$ $16.36$
DFS S-DDQ DDM $23.04$ $18.84$ $19.56$ $17.48$
DFS SC-DDQ DME $35.16$ $36.54$ $17.28$ $26.16$
DFS SC-DDQ DEE $35.16$ $16.68$ $13.61$ $24.27$
DFS SC-DDQ DDM $35.16$ $41.93$ $35.12$ $31.40$
5.1 The Effectiveness of the Proposed SC-DDQ and Its Variants
According to the experimental results, SC-DDQ and S-DDQ achieved outstanding performance compared with classic dialog agents, such as DQN and DDQ. As illustrated in Table 4, schedule EMD yielded the highest success rate on SC-DDQ (0.92) at the final training stage, whereas schedules DEE and DDM provided the best performance for S-DDQ (0.94), outperforming DQN (0.74) and DDQ (0.86) by a large margin. Fig. 8 shows the relationship between the success rate and the number of turns. This result clearly shows that a high success rate and small number of turns were achieved with better training strategies. In particular, good results were obtained with DDQ using the difficult-first strategy (DFS) without curiosity (S-DDQ), or DDQ using the easy-first strategy (EFS) with curiosity (SC-DDQ). S-DDQ with DFS yielded the best success rate, whereas SC-DDQ with EFS achieved the smallest number of turns.
Figure 7: Success rates and average turns.
Figure 7: Success rates and average turns.
Figure 8: Success rates at each stage.
5.2 The Effectiveness of Two Opposite Training Strategies
Both typical CL and its reverse version (difficult-first strategy, DFS) benefit from the dialog policy without or with curiosity, respectively. Fig. 8 shows the success rates for each training stage. The results for the same condition were averaged. For example, the curiosity_DFS line is the average of the SC-DDQ with DME, DDE, and DDM schedules. Overall, schedules designed by the typical CL (EFS) work better on agents with the curiosity policy, and DFS is beneficial for those without the curiosity policy. Because the training of agents started with RBS, where the same rule-based policy designed for dealing with easy tasks was adopted, agents exposed to easy tasks initially showed a rapid performance boost, and the corresponding success rate reached 1.00. As the difficulty level of subsequent tasks increased, success rates decreased. Specifically, when the model without the curiosity policy attempted to grasp difficult samples after exposure to easy ones, its success rate decreased drastically, leading to poor training results in the final phase. For the model with the curiosity policy, following EFS, its success rate decreased slightly in the middle of training; however, it picked up considerably towards the end of training thanks to the encouragement of state exploration from the curiosity model.
User goal:
"request_slots": {
"ticket": "UNK"
"diaact": "request",
"inform slots": {
"city": "Birmingham",
"number of people": "4",
"theater": "Carmike summit 16",
"state": "al",
"start time": "around 6pm",
t "date": "today",
"movie name": "deadpool"
Dialog conducted by rule-based agent:
Turn 0 usr:
Can I get 4 tickets for deadpool?
Turn 1 sys:
Which city would you like?
Turn 2 usr:
I want to watch at Birmingham.
Turn 3 sys:
Carmike summit 16 is available.
Turn 4 usr:
I want to watch at al.
Turn 5 sys:
What time would you like to see it?
Turn 6 usr:
I want to watch at around 6pm.
Turn 7 sys:
What date would you like to watch it?
Turn 8 usr:
I want to set it up today.
Turn 9 sys:
Great, I was able to purchase 4 tickets to deadpool for today at around 6pm at Carmike summit 16 in Birmingham, al.
Turn 10 usr:
Thank you.
(a) An easy user goal
User goal:
"request_slots": {
"date": "UNK",
"theater": "UNK",
"start time": "UNK",
"ticket": "UNK"
"diaact": "request",
"inform slots": {
"city": "Portland",
"number of people": "4",
"state": "Oregon",
"movie name": "star wars"
Dialog conducted by rule-based agent:
Turn 0 usr:
When is star wars playing in Oregon?
Turn 1 sys:
Portland is available.
Turn 2 usr:
What date is available?
Turn 3 sys:
12:05pm is available.
Turn 4 usr:
What date is available?
Turn 5 sys:
Friday is available.
Turn 6 usr:
Which theater is available?
Turn 7 sys:
Regal Lloyd center 10 & IMAX is available.
Turn 8 usr:
I want 4 tickets please!
Turn 9 sys:
Great, I was able to purchase 4 tickets to star wars for Friday at 12:05pm at regal Lloyd center 10 & IMAX in Portland, Oregon.
Turn 10 usr:
Thank you.
(b) A difficult user goal
Figure 9: Dialogs with different user goals
When the agent first focused on difficult samples, the success rate at the early stage was lower than that of EFS. However, as the difficulty of the task decreased and the policy was improved, the success rates increased. In the final stage, the success rates of both agents optimized according to DFS decreased, except for schedule DDM.
(a) Stage 1: Easy tasks
(b) Stage 2: Easy tasks
(c) Stage 3: Difficult tasks
(d) Stage 4: All tasks
Figure 10: Distributions of actions sampled by S-DDQ EED.
5.3 Observation of Entropy of Action Sampling
From the results shown in the previous section, we obtained a high success rate under two conditions: S-DDQ with DFS and SC-DDQ with EFS. To investigate why these conditions gave good results and other conditions did not, we introduced the entropy of sampled actions, which is used to depict the behavioral characteristics of agents. Agent sampling action preferences and entropy calculations are explained in the following paragraph. When training at different task levels, the agent prefers to sample different actions. when facing easy tasks, as shown in Fig. 9 (a), the agent tends to ask more questions instead of providing information. When facing difficult ones, as shown in Fig. 9 (b), the actions sampled by the agent tend to diversify. In other words, the agent attempts to provide appropriate information to answer the user query instead of simply asking the user questions. Furthermore, the order in which the agent meets tasks of varying difficulty affects action sampling. Therefore, to explore the effect of training schedules on agent performance, we counted the actions employed by the agent in each training phase and calculated the frequency with which each action was sampled. For example, Fig. 10 (a), (b), (c), and (d) show the frequency of actions sampled by the agent without curiosity equipped with schedule EED. On schedule EED, the agent is sequentially trained with easy, easy, and difficult tasks for 70 epochs per training stage. Finally, it is trained with all tasks for 90 epochs. The vertical coordinate is the frequency and the horizontal coordinate represents the 29 possible actions. The entropy of the distribution of sampled actions is formulated as
$H=-\sum_{i=0}^{28}{P\left(a_{i}\right)\log{P(a_{i})}}$ (2)
Here, $a_{i}$ represents feasible actions. $P(a_{i})$ is the probability of choosing action $a_{i}$ during the current training stage. Entropy represents the degree of uniformity in sampling. A larger entropy implies that the agent samples actions more evenly and attempts to perform various actions rather than being trapped in a few pseudo-optimal actions. Table 6 lists the entropy values at each stage for all training conditions. Roughly speaking, in the S-DDQ with EFS conditions, the entropy values were small in the first stage and gradually increased in the following stages. The entropy values of S-DDQ with DFS showed an opposite trend. In addition, SC-DDQ with EFS showed large entropy in the first stage, whereas that with DFS showed a random up-down trend.
Table 6: Entropy values of each stage for all training conditions.
Strategy Method Schedule S1 S2 S3 S4
Random DDQ - 3.98 4.16 3.76 3.63
Random C-DDQ - 4.15 3.61 3.55 3.6
EFS S-DDQ EMD 3.53 3.58 3.97 3.93
EFS S-DDQ EDD 3.53 3.68 3.87 3.55
EFS S-DDQ EED 3.53 2.88 3.96 3.84
EFS SC-DDQ EMD 4.04 3.77 4 3.76
EFS SC-DDQ EDD 4.04 4.08 3.94 3.96
EFS SC-DDQ EED 4.04 3.29 4.05 3.87
DFS S-DDQ DME 4.17 4.06 3.67 3.77
DFS S-DDQ DEE 4.17 3.43 3.32 3.87
DFS S-DDQ DDM 4.17 3.98 3.69 3.57
DFS SC-DDQ DME 4.05 4.1 3.68 4.42
DFS SC-DDQ DEE 4.05 3.45 3.37 4.23
DFS SC-DDQ DDM 4.05 4.19 4.28 4.08
6 Discussion
Unlike traditional CL, in the present study, we observed that EFS and DFS had different effects on task completion ability in different task-oriented dialog models. DFS is beneficial for S-DDQ, whereas the classic CL is helpful for SC-DDQ. These two combinations share high entropy in the first stage and low entropy in the last stage, where sufficient action exploration leads to better performance. Contrary to intuition, employing a curiosity model designed to explore states is not always effective for exploring actions. To investigate the effects of the final success rate, we calculated Pearson’s correlation coefficient between the entropy values at each stage and the final success rate. The results are shown in Fig. 11. This result shows that entropy in the first stage positively correlates with the success rate, whereas entropy in the final stage is negatively correlated with the success rate. This means that an agent with a high success rate samples the actions more randomly first, and the sampling diversity converges in the last stage. If curiosity is not employed, training with easy tasks leads to a greater sampling bias and smaller entropy. This situation can be avoided using difficult tasks or curiosity. However, if we use difficult tasks at an early stage when using curiosity, the action sampling model does not converge and the success rate does not improve.
Figure 11: Correlation between entropy values at each stage and the success rate at the final stage.
7 Limitations and Future works
This research has several limitations. First, the training schedules offered in this research were manually designed. Such schedules lack flexibility, and once designed, changes can not be made during the training process. Second, the curiosity reward cannot be adaptively adjusted for the time being. It participates in agent optimization with the same weight from beginning to end. However, dynamic adjustment of curiosity is necessary for its application in different scenarios. Finally, in the early training stage, the curiosity reward corresponding to each feasible action may deviate from its actual value. Compared with ICM [11], our curiosity model outputs the curiosity reward directly before updating to the next state, which allows the agent to obtain curiosity rewards before executing the action. Because the agent does not actually execute feasible actions individually and observe the next state, it does not obtain the real next state before acquiring the curiosity reward, and the curiosity reward output by the curiosity model is a simulation. By analyzing the entropy of action sampling, we found that the dialog system exhibits better performance when the entropy tends to decrease. In other words, we advocate that the RL-based dialog agent could benefit from encouraging exploration in the early training stage and then gradually decreasing the level of exploration. Based on our findings, we plan to add weight to the curiosity reward to make curiosity-driven exploration more adaptive. Further work on the gradual discarding of curiosity is worth conducting, because this is a generic promising path for other RL problems.
8 Conclusion
This paper presents a new framework, Scheduled Curiosity-DDQ (SC-DDQ), for task-oriented dialog policy learning. With the introduction of a curiosity model and scheduled training strategy, SC-DDQ outperforms previous classic dialog agents, such as DDQ and DQN. Moreover, variants of SC-DDQ were conducted to verify the effectiveness and influence of a typical CL and its reverse version. Based on the experimental results, the difficult-first strategy benefits S-DDQ, whereas the easy-first strategy is preferable for SC-DDQ. To explore the impact of task difficulty on the dialog agent policy, we calculated entropy. We found common trends of entropy, where the agent tried various actions randomly in the first stage, and converged in the last stage. In the future, we plan to control the curiosity-driven exploration during the training process based on entropy trends. Furthermore, we believe that this curriculum framework can be applied to improve other RL-based NLP models.

### v2

@@
 Figure 1: The framework of DDQ
-Figure 1: The framework of DDQ
 Figure 2: The structure of world model
@@
 Dialog conducted by rule-based agent:
-Turn 0 usr:
-I want tickets for race at 10:00 pm.
-Turn 1 sys:
-What movie are you interested in?
-Turn 2 usr:
-I want to watch race.
-Turn 3 sys:
-What time would you like to see it?
-Turn 4 usr:
-I want to watch at 10:00 pm.
-Turn 5 sys:
-Which city would you like?
-Turn 6 usr:
-I want to watch at Seattle.
-Turn 7 sys:
-What date would you like to watch it?
-Turn 8 usr:
-I want to set it up tomorrow
-Turn 9 sys:
-Which theater would you like?
-Turn 10 usr:
-I want to watch at amc pacific place 11 theater.
-Turn 11 sys:
-How many tickets do you need?
-Turn 12 usr:
-I want 2 tickets please!
-Turn 13 sys:
-Great - I was able to purchase 2 tickets for you to see race tomorrow at amc pacific place 11 theater in Seattle at 10:00 pm.
-Turn 14 usr:
-Thank you.
-Turn 15 sys:
-Thank you.
+Turn 0 usr: I want tickets for race at 10:00 pm.
+Turn 1 sys: What movie are you interested in?
+Turn 2 usr: I want to watch race.
+Turn 3 sys: What time would you like to see it?
+Turn 4 usr: I want to watch at 10:00 pm.
+Turn 5 sys: Which city would you like?
+Turn 6 usr: I want to watch at Seattle.
+Turn 7 sys: What date would you like to watch it?
+Turn 8 usr: I want to set it up tomorrow
+Turn 9 sys: Which theater would you like?
+Turn 10 usr: I want to watch at amc pacific place 11 theater.
+Turn 11 sys: How many tickets do you need?
+Turn 12 usr: I want 2 tickets please!
+Turn 13 sys: Great - I was able to purchase 2 tickets for you to see race tomorrow at amc pacific place 11 theater in Seattle at 10:00 pm.
+Turn 14 usr: Thank you.
+Turn 15 sys: Thank you.
 Figure 4: A success task completed by rule-based agent
 Table 2: The data annotation schema
-Intent
-request, inform, deny, confirm_question,
+Intent request, inform, deny, confirm_question,
 confirm_answer, greeting, closing, not_sure,
 multiple_choice, thanks, welcome
-Slot
-city, closing, date, distanceconstraints,
+Slot city, closing, date, distanceconstraints,
 greeting, moviename, numberofpeople, price,
@@
 theater_chain, ticket, video_format, zip
+In this research, we designed an offline classifier focusing on booking movie tickets, identifying user goals into three complexity levels: easy, middle, and difficult.
+A user goal comprises two parts: request and inform slots.
+Inform slots are constraints known or determined by the user.
+The request slots are unknown to the user and must be obtained from the responses of the agent.
+Table 2 lists all slots.
+Consider Fig.
+5 (b) as an example.
+This user goal contains three request slots: ticket, theater, and start time, and three inform slots: number of people, date, and movie name.
+This goal reveals that the user would like to buy three tickets for Zootopia tonight, but he/she has yet to decide which theater to go to and at a specific time.
+During this episode of dialog, the agent should provide information about a suitable theater and start time to satisfy the user’s constraints.
+In addition, the ticket is a default slot that always appears in the request slots for user goals.
+Before launching the iterative scheduled SC-DDQ, the system is opened up with a warm starting phase, where a rule-based dialog policy [7] is used to interact with users and generate experiences.
+As the dialog strategy of the rule-based agent requires the movie name, start time, city, date, theater, and number of people in order, a user goal with a limited number of request slots is a breeze for an agent to achieve, as a successful example shown in Fig.
+4.
+Therefore, this research applies the number of request slots as the measurement of the difficulty of training tasks, rather than the length of dialog, word rarity, and the total number of requests and inform slots [34, 13, 35, 36, 18].
+The goal classification was performed offline.
+In this study, we classified user goals by the number of request slots, as listed in lines 2-12 of Algorithm 1.
+Fig.
+5 shows additional examples of the user goals at different levels.
+As shown in Fig.
+5(a), goals with only one request slot are classified as easy and stored in the easy goal buffer $G_{\textnormal{easy}}$.
+Goals with two or three slots are then classified as middle and stored in the middle goal buffer $G_{\textnormal{middle}}$.
+Finally, goals with four or five slots are classified as difficult and stored in the difficult goal buffer $G_{\textnormal{difficult}}$.
+At the beginning of the dialog, the user samples a goal from the corresponding set.
+In this research, schedules designed on two opposite training strategies are conducted, which are Easy-Middle-Difficult (EMD), Easy-Difficult-Difficult (EDD), Easy-Easy-Difficult-All (EED), Difficult-Middle-Easy (DME), Difficult-Easy-Easy (DEE) and Difficult-Difficult-Middle (DDM).
 […]
@@
 27       end while
+28
 29      Update $\theta_{Q}$ using real experiences from $D^{u}$ via minibatch SGD;
@@
 33       Update $\theta_{C}$ using real and simulated experiences via minibatch SGD;
-33
+34
 35 end for
@@
 Figure 7: Success rates and average turns.
-Figure 7: Success rates and average turns.
 Figure 8: Success rates at each stage.
@@
 User goal:
-"request_slots": {
-"ticket": "UNK"
-"diaact": "request",
-"inform slots": {
-"city": "Birmingham",
-"number of people": "4",
-"theater": "Carmike summit 16",
-"state": "al",
-"start time": "around 6pm",
-t "date": "today",
-"movie name": "deadpool"
+{ "request_slots": { "ticket": "UNK" }, "diaact": "request", "inform slots": { "city": "Birmingham", "number of people": "4", "theater": "Carmike summit 16", "state": "al", "start time": "around 6pm", t "date": "today", "movie name": "deadpool" } }
 Dialog conducted by rule-based agent:
-Turn 0 usr:
-Can I get 4 tickets for deadpool?
-Turn 1 sys:
-Which city would you like?
-Turn 2 usr:
-I want to watch at Birmingham.
-Turn 3 sys:
-Carmike summit 16 is available.
-Turn 4 usr:
-I want to watch at al.
-Turn 5 sys:
-What time would you like to see it?
-Turn 6 usr:
-I want to watch at around 6pm.
-Turn 7 sys:
-What date would you like to watch it?
-Turn 8 usr:
-I want to set it up today.
-Turn 9 sys:
-Great, I was able to purchase 4 tickets to deadpool for today at around 6pm at Carmike summit 16 in Birmingham, al.
-Turn 10 usr:
-Thank you.
+Turn 0 usr: Can I get 4 tickets for deadpool?
+Turn 1 sys: Which city would you like?
+Turn 2 usr: I want to watch at Birmingham.
+Turn 3 sys: Carmike summit 16 is available.
+Turn 4 usr: I want to watch at al.
+Turn 5 sys: What time would you like to see it?
+Turn 6 usr: I want to watch at around 6pm.
+Turn 7 sys: What date would you like to watch it?
+Turn 8 usr: I want to set it up today.
+Turn 9 sys: Great, I was able to purchase 4 tickets to deadpool for today at around 6pm at Carmike summit 16 in Birmingham, al.
+Turn 10 usr: Thank you.
 (a) An easy user goal
@@
 Dialog conducted by rule-based agent:
-Turn 0 usr:
-When is star wars playing in Oregon?
-Turn 1 sys:
-Portland is available.
-Turn 2 usr:
-What date is available?
-Turn 3 sys:
-12:05pm is available.
-Turn 4 usr:
-What date is available?
-Turn 5 sys:
-Friday is available.
-Turn 6 usr:
-Which theater is available?
-Turn 7 sys:
-Regal Lloyd center 10 & IMAX is available.
-Turn 8 usr:
-I want 4 tickets please!
-Turn 9 sys:
-Great, I was able to purchase 4 tickets to star wars for Friday at 12:05pm at regal Lloyd center 10 & IMAX in Portland, Oregon.
-Turn 10 usr:
-Thank you.
+Turn 0 usr: When is star wars playing in Oregon?
+Turn 1 sys: Portland is available.
+Turn 2 usr: What date is available?
+Turn 3 sys: 12:05pm is available.
+Turn 4 usr: What date is available?
+Turn 5 sys: Friday is available.
+Turn 6 usr: Which theater is available?
+Turn 7 sys: Regal Lloyd center 10 & IMAX is available.
+Turn 8 usr: I want 4 tickets please!
+Turn 9 sys: Great, I was able to purchase 4 tickets to star wars for Friday at 12:05pm at regal Lloyd center 10 & IMAX in Portland, Oregon.
+Turn 10 usr: Thank you.
 (b) A difficult user goal